# DXAI: Explaining Classification by Image Decomposition

### v1

@@ 1 Introduction
Understanding classification reasoning of neural networks is of paramount importance. It can lead to better understanding of the classificaction process, help in debugging and validation and also serve as an additional informative output for the user at inference. Thus, the research on the topic of explainable artificial intelligence (XAI) is extensive [11, 18, 2, 13].
Figure 1: What makes a cat a Cat? Explained through decomposition. Examples of class-distinct, $\psi_{Distinct}$, and class-agnostic, $\psi_{Agnostic}$, parts of Animals dataset [8], where the input $x=\psi_{Agnostic}+\psi_{Distinct}$. The distinct part gives high resolution, dense, color and texture information explaining the classification decision. This is in contrast to standard XAI (bottom two rows).
@@
In such scenarios, a heatmap is much less informative. Depending on the algorithm, it is either too focused on just a small portion of highly dominant features or it shows uniform large areas spanning most of the image. See some examples in Fig. 3. In both cases, we lack a clear constructive explanation for the network’s decision.
Figure 2: Three visualizations using a heatmap. Generated by Grad-CAM [29], highlighting cars in an aerial image (left).
Figure 3: Our method vs. heatmaps, illustrating three scenarios mentioned in Section 1: 1) Many details spread across large portions of the image are helpful for accurate classification (top row), heatmaps show only partial relevant information. 2) Distinguishing between types of peppers that differ mainly by color (2nd row), heatmaps struggle to provide color data. 3) Detecting additive statistical disturbance (a class of clean images and a class of images with noise). Here as well, heatmaps face difficulties explaining the reason for classification.
To alleviate the above problems, we propose a different way of computing and visualizing explanations of the classification. It does not use heatmaps and is based on decomposition. Image decomposition is a long standing paradigm in classical image processing. It is based on the notion that images (and signals in general) can often be well decomposed into meaningful additive parts. Given a 1D signal $x$ of length $N$ and some orthonormal basis of $N$ elements $\{\phi_{i}\}$, $i=1,..\,N$, one can express $x$ as
@@
There are many other studies attempting to decompose an image into meaningful additive parts. For example, image decomposition into structure and texture $x=\psi_{Structure}+\psi_{Texture}(+e)$ (with possible noise or small error) [4, 3, 12].
@@ 2 Related Work
    Single channel. Typically, these algorithms produce grayscale images representing the importance of each pixel, lacking color distinction (multi-channel data in general). This can lead to inadequate explanations, as seen, for example, in our peppers dataset experiment (Fig. 11).
@@ 3.1 The DXAI problem
We first introduce the basic setting and notations. We assume a multiclass supervised learning setting. Let $X$ be the space of input images and $x\in X$ be an input image belonging to one of $c$ classes. Let $y\in Y$ be a class label $y\in\{0,c-1\}$. The training set consists of $M$ pairs $(x_{i},y_{i})\in X\times Y$, $i=1,..\,,M$. We denote by $x^{y}$ an image $x$ belonging to class $y$. We are given an image classifier $C$, which provides a class probability vector $p$. We would like to extract the relevant features, and only them, which contribute to the classification decision by the network. An additive contribution, as in Eq. (1), is assumed. We want to have a reference image, which is neutral in terms of classification (for a given classifier), using the following definition:
@@
In words, for a given image $x$, we find the closest image (with respect to $d$) which is neutral, in terms of class (agnostic part). The difference between $x$ and this neutral image explains the probability vector $p(x)$ and the reason, according to the classifier, why it deviates from neutrality. We do not assume a unique solution. The pair $\{\psi_{Agnostic},\psi_{Distinct}\}$ provide detailed and dense class explanation. We note that pure agnostic images are hard to obtain. A more relaxed requirement is to have $|p_{i}(\psi)-\frac{1}{c}|\leq\varepsilon$, $0<\varepsilon\ll\frac{1}{c}$ in Def. 1. In any case, problem (2) is very hard to solve. We give below a suggested approximation. Alternative methods may be further investigated.
@@ 3.3 Training
Orthogonality loss. We also incorporate a loss function to promote low correlation between branches in the feature space. This loss is defined as:
$L_{orth}=\sum_{l=1}^{L}\sum_{i\neq j}(\textrm{corr}(F^{l}_{i},F^{l}_{j})-I)^{2},$ (9)
where corr is a correlation matrix, $F^{l}_{i}$ are the features of the $i$’th generator at layer $l$ and $I$ is the identity matrix. Although this loss is not mandatory, we noticed it reduces overall correlation of $\psi$’s and emphasizes unique characteristics in each component.
@@
Figure 8: BraTS [25]: Tumors in MRI images, class-distinct and class-agnostic parts. The dataset included two classes, with and without tumors.
Ours LRP IG G-SH II GGC GC
0.125 0.156 0.171 0.172 0.177 0.179 0.192
Table 2: AUC comparison of the Faithfulness index on the BraTS dataset among various XAI methods. Ours is compared to LRP, Integrated gradients, Gradient SHAP, Internal influence, Guided Grad-Cam and Grad-Cam. AUC values were calculated based on the deletion of pixels, with up to 20% removed. When our CD part may be interpreted as importance (by pixel amplitude) one can validate our accurate XAI mapping also by standard evaluation methods.
DOTA (cars) AFHQ Celeb A
AUC according to: ResNet18 Simple ResNet18 Simple ResNet18 Simple
Trained by ResNet18 0.793 0.903 0.737 0.759 0.616 0.707
Trained by Simple-Classifier 0.914 0.729 0.865 0.623 0.69 0.609
Table 3: Matching explanation to classifier. At each row DXAI training is performed using a different classifier. AUC (similiar to Tab. 1) is shown, computed using the two classifiers, on three different data sets. AUC is smaller (better) when the classifier in training and the one used to compute the AUC match. This shows the relevance of the DXAI explanations to the classifier.
Figure 9: Cars (DOTA [34]): Class-distinct and Class-agnostic parts. The dataset included two classes, images with and without cars.
Figure 10: CelebA [15]: Class-distinct and Class-agnostic parts. The dataset included male and female classes.
Figure 11: Peppers [31]: Class-distinct and Class-agnostic parts. The dataset included four classes of different colored peppers.
4 Experiments
We show here various experiments examining the validity of our algorithm and the applicability of using the proposed method to infer useful classification explanations for various diverse data sets. Unless noted otherwise, we use ResNet18 [16] as our image classifier. The classification problem is solved sufficiently well in all cases with about 95% or higher test accuracy.
$x_{\beta}=x-\beta\cdot\psi_{Distinct},$ (10)
@@ 4 Experiments
Our algorithm does not provide importance ranking of each pixels, with respect to its contribution for the classification. This is one of its limitations. Thus we cannot use the standard way to evaluate XAI accuracy by removing pixels gradually based on importance, as done e.g. in [28]. In some problems, however, a simple ranking can be inferred by our algorithm. In a binary classification problem, when one class is predominantly decided based on the existence of certain lighter pixels, we can use the amplitude of pixels in $\psi_{Distinct}$ as a reasonable importance ranking. In this case standard XAI evaluation can be made. The BraTS dataset [25] is such a case. It contains MRI scans of the human brain used for brain tumor segmentation research. The dataset includes scans from patients with brain tumors, along with expert annotations for tumor regions. In many cases, bright regions indicate evidence of tumors. We divided the dataset into two classes: images containing tumors and images that do not. Some example results are shown in Fig. 8. Table 2 shows the standard AUC evaluation on this set (AUC of accuracy vs. number of pixels removed, ordered by importance). We see our algorithm behaves favorably both qualitatively and quantitatively.
@@
In Figs. 9, 10 and 11 we show additional examples for various datasets. In general, our method is especially effective in cases where the differences between the classes have an additive nature. This occurs in situations where the differences involve textures (head- and facial-hair in Fig. 10), colors (Fig. 11), and the presence or absence of details, such as cars (Fig. 9), red lips or heavy eyebrows (Fig. 10).
@@ 5 Discussion and Conclusion
We propose an alternative way to analyze and to visualize the reasons for classification by neural networks. It is based on decomposing the image into a part which does not contribute to the classification and to one which holds the class-related cues. This approach may not be ideal for all applications and has several limitations and drawbacks, compared to standard XAI methods: 1. The method requires training, for a specific training set and classifier (training is slow, inference is fast); 2. There is no natural ranking of the significance of pixels in the image. The amplitude of pixels in the class-distinct part can serve as a good approximation; 3. Our implementation uses GANs. We note that image realism is of secondary significance for XAI. Our concept does not rely on a GAN architecture and improvements may be achieved by diffusion-type generative models, such as [17, 36, 33, 7].
@@
Appendix A Detailed description of the method
A.1 Training
$L_{KLD}=KLD\Bigl{(}D(x^{y}),C(x^{y})\Bigl{)},$ (12)
where $KLD$ is the Kullback-Leibler distance.
$d(u,v)=\|u-v\|_{L^{1}}+\|u-v\|_{L^{2}},$
with a reconstruction loss of the form
$L_{rec}=d(x,\hat{x}^{y})+d(x,\hat{x}^{\tilde{y}}).$
The proposed loss function is
$L_{dis-rec}=d(x\odot\mathbb{I},\hat{x}^{y}\odot\mathbb{I}),$ (13)
where
and $\odot$ denotes element-wise product.
The full objective loss is:
$L_{Total}=\lambda_{adv}L_{adv}+\lambda_{KLD}L_{KLD}+\lambda_{cf}L_{class-fake}$ (14)
$+\lambda_{rec}L_{rec}+\lambda_{dr}L_{dis-rec}+\lambda_{o}L_{orth}.$
In our experiments we used the following weights:
$\lambda_{adv}=2,\lambda_{KLD}=1,\lambda_{cf}=2,$ (15)
$\lambda_{rec}=4,\lambda_{dr}=4,\lambda_{o}=1.$
A.2 Inference Stage
@@
Figure 13: Noise distinction. Celeb-A where one class contains clean images and the other class contains noisy images, contaminated by additive white Gaussian noise (AWGN). In this case, standard XAI algorithms yield partial and inaccurate explanation of the class distinction. Our algorithm provides a correct explanation distincting the two classes by isolating well the noise.
A.3 Datasets
No modifications were made to the remaining datasets.
A.4 Classifiers
In our experiments three different classifiers were employed for the classification block $C$: a ResNet18-type classifier, a simple classifier described in Tab. 6, and our discriminator $D$, which possesses classification capabilities, as the output is a vector of length $c$, indicating class probabilities (details of the architecture are in Tab. 7). The test accuracy of each classifier is given in Tab. 5. We note that the discriminator exhibits the best classification performance. Thus, we selected it in the main paper to show the explanation images by the various DXAI and XAI algorithms. Note that we found a mistake in the explanation text of the paper, which will be corrected in the final version. We conducted experiments also on ResNet18 and on the Simple classifier, showing consistently similar trends. We provide later some examples of ResNet18 results on the same datasets.
@@ A.4 Classifiers
The intentionally designed simple classifier, with lower classification ability, provides distinct classification results from a well-known classifier like ResNet18. This intentional dissimilarity allows for meaningful comparisons of their results, as detailed in the main paper.
Figure 14: AFHQ ablation - need of more than 2 branches. The run of the experiment was stopped after 150K iterations, after the tendency was understood.
Figure 15: CelebA ablation - need of more than 2 branches. The run of the experiment was stopped after 150K iterations, after the tendency was understood.
Appendix B Ablation
@@
In the ablation study, we investigated two key factors. Firstly, we examined the impact of the number of branches. While decomposing the classified image into several images, we primarily focus on solving only two images—distinct and agnostic. One might question why not use only two branches for the solution, since two branches can be trained to achieve a similar solution.
@@
We show that when branches are used for the agnostic part results are better both in terms of reconstruction and of the generators’ ability to produce images that explain the classifier. For instance, as demonstrated in Figs. 14 and 15, PSNR decreases when using only two branches. Additionally, the loss $L_{class-fake}$, representing the generators’ ability to produce meaningful images of a specific class according to the given classifier, is higher for the entire training with two branches. In other words, the classifier interprets the images less accurately as the desired class, making the images less reliable. The additional generators provide better generation capacity and can be trained in an easier and more stable manner.
@@
Appendix C More results

### v2

@@ 1 Introduction
Figure 1: Our method vs. heatmaps, illustrating three scenarios where heatmaps are less informative: 1) Many details spread across large portions of the image are helpful for accurate classification (top row), heatmaps show only partial relevant information. 2) Distinguishing between types of peppers that differ mainly by color (2nd row). 3) Detecting additive statistical disturbance (a class of clean images and a class of images with noise). Since the contribution is global – heatmaps face difficulties explaining the reason for classification.
Understanding classification reasoning of neural networks is of paramount importance. It can lead to better understanding of the classificaction process, help in debugging and validation and also serve as an additional informative output for the user at inference. Hence, the research on the topic of explainable artificial intelligence (XAI) is extensive [11, 17, 2, 13].
@@
In such scenarios, a heatmap is much less informative. Depending on the algorithm, it is either too focused on just a small portion of highly dominant features or it shows uniform large areas spanning most of the image. See some examples in Fig. 1. In both cases, we lack a clear constructive explanation for the network’s decision. Furthermore, the basis for using explanations through heatmaps implicitly assumes that some pixels are crucial for the solution while others are either unimportant or only partially important. We argue that there are scenarios where each pixel contains both class identity information and neutral data, necessitating the use of a different method.
Figure 2: Three visualizations using a heatmap. Highlighting cars in an aerial image (left). Generated by Grad-CAM [28].
Our approach assumes a membership logic, such that each region is potentially a superposition of image features common to many classes and ones which are class-specific. Image decomposition is a long standing paradigm in classical image processing. It is based on the notion that images (and signals in general) can often be well decomposed into meaningful additive parts. Given a 1D signal $x$ of length $N$ and some orthonormal basis of $N$ elements $\{\phi_{i}\}$, $i=1,..\,N$, one can express $x$ as
@@
There are many other studies attempting to decompose an image into meaningful additive parts. For example, image decomposition into structure and texture $x=\psi_{Structure}+\psi_{Texture}(+e)$ (with possible noise or small error) [34, 4, 3, 12]. See example in Fig. 3(a).
@@
(a) Additive decompositions. The top image illustrates texture-structure decomposition [34]. Our explanation is similarly additive, for peppers classification we implicitly obtain a structure-color decomposition.
/ Refer to caption
(b) Peppers [30]: Class-distinct and Class-agnostic parts. The dataset includes four classes of peppers of different color.
Figure 3: An image can often be decomposed into meaningful additive parts, such as in classical structure-texture decomposition (top-left). This example shows how it allows to explain well color-based reasoning.
@@ 2 Related Work
    Single channel. Typically, these algorithms produce grayscale images representing the importance of each pixel, lacking color distinction (multi-channel data in general). This can lead to inadequate explanations, as seen, for example, in our peppers dataset experiment (Fig. 3(b)).
@@ 3.1 The DXAI problem
We first introduce the basic setting and notations. We assume a multiclass supervised learning setting. Let $X$ be the space of input images and $x\in X$ be an input image belonging to one of $c$ classes. Let $y\in Y$ be a class label $y\in\{0,c-1\}$. The training set consists of $M$ pairs $(x_{i},y_{i})\in X\times Y$, $i=1,..\,,M$. Let $x^{y}$ denote an image $x$ belonging to class $y$. An image classifier $C$ is given, which provides a class probability vector $p$.
We would like to extract the relevant features, and only them, which contribute to the classification decision by the network. An additive contribution, as in Eq. (1), is assumed. We want to have a reference image, which is neutral in terms of classification (for a given classifier), using the following definition:
@@
In words, for a given image $x$, we find the closest image (with respect to $d$) which is neutral, in terms of class (agnostic part). The difference between $x$ and this neutral image explains the probability vector $p(x)$ and the reason, according to the classifier, why it deviates from neutrality. We do not assume a unique solution. The pair $\{\psi_{Agnostic},\psi_{Distinct}\}$ provide detailed and dense class explanation. Obtaining an approximate agnostic part may be computed for each image directly using optimization techniques. For example, by minimizing a loss that requires minimal distance in terms of KL divergence between the distribution generated by the classifier $C$ and a uniform distribution. However, the distinction component is not semantically viable, it is mostly based on out-of-distribution features (which resemble noise, see the example in Appendix 0.E). Obtaining a solution that captures semantic characteristics is more challenging. We propose to solve this using generative models.
@@ 3.3 Training
4 Experiments
We show here various experiments examining the validity of our algorithm and the applicability of using the proposed method to infer useful classification explanations for various diverse data sets. During the experiments we used three different classifiers. Details about their architecture appear in the supplementary.
In Figs. 7, 8 and 3(b) we show qualitative examples for various datasets. In general, our method is especially effective in cases where the differences between the classes have an additive nature. This occurs in situations where the differences involve textures (head- and facial-hair in Fig. 7(a)), colors (Fig. 3(b)), and the presence or absence of details, such as cars (Fig. 7(c)), tumors (Fig. 7(d)) red lips or heavy eyebrows (Fig. 7(a)).
(a) CelebA [15]: The dataset included male and female classes.
(b) AFHQ [8]: The dataset includes animals faces.
(c) Cars (DOTA [33]): Aerial images. The dataset included two classes, images with and without cars.
(d) BraTS [24]: Tumors in MRI images. The dataset included two classes, with and without tumors.
Figure 7: Comparison of class-distinct and class-agnostic parts with heatmaps from other methods.
Figure 8: Noise distinction. One class contains clean images (face images from Celeb-A) and the other class contains noisy images (additive white Gaussian noise). In this case, standard XAI algorithms yield partial and inaccurate explanation of the class distinction. Our algorithm gives a correct explanation, providing the noise part as the reason for classification.
@@ 4 Experiments
Our algorithm does not provide importance ranking of each pixels, with respect to its contribution for the classification. This is one of its limitations. Thus we cannot use the standard way to evaluate XAI accuracy by removing pixels gradually based on importance, as done e.g. in [27]. In some problems, however, a simple ranking can be inferred by our algorithm. In a binary classification problem, when one class is predominantly decided based on the existence of certain lighter pixels, we can use the amplitude of pixels in $\psi_{Distinct}$ as a reasonable importance ranking. In this case standard XAI evaluation can be made. The BraTS dataset [24] is such a case. It contains MRI scans of the human brain used for brain tumor segmentation research. The dataset includes scans from patients with brain tumors, along with expert annotations for tumor regions. In many cases, bright regions indicate evidence of tumors. We divided the dataset into two classes: images containing tumors and images that do not. Some example results are shown in Fig. 7(d). Table 2 shows the standard AUC evaluation on this set (AUC of accuracy vs. number of pixels removed, ordered by importance). We see our algorithm behaves favorably both qualitatively and quantitatively.
Table 2: AUC comparison of the Faithfulness index on the BraTS dataset among various XAI methods. Ours is compared to LRP, Integrated gradients, Gradient SHAP, Internal influence, Guided Grad-Cam, Grad-Cam and random noise. AUC values were calculated based on the deletion of pixels, with up to 20% removed. When our CD part may be interpreted as importance (by pixel amplitude) one can validate our accurate XAI mapping also by standard evaluation methods.
Ours LRP [5] IG [29] G-SH [18] I-Inf [20] GGC [28] GC [28] RAND
0.125 0.156 0.171 0.172 0.177 0.179 0.192 0.195
Table 3: Matching explanation to classifier. At each row DXAI training is performed using a different classifier. AUC (similiar to Tab. 1) is shown, computed using the two classifiers, on four different data sets. AUC is smaller (better) when the classifier in training and the one used to compute the AUC match. This shows the relevance of the DXAI explanations to the classifier.
BraTS DOTA (cars)
AUC according to: ResNet18 Simple ResNet18 Simple
Trained by ResNet18 0.49 0.787 0.793 0.903
Trained by Simple-Classifier 0.629 0.562 0.914 0.729
AFHQ Celeb A
AUC according to: ResNet18 Simple ResNet18 Simple
Trained by ResNet18 0.61 0.871 0.606 0.807
Trained by Simple-Classifier 0.917 0.546 0.809 0.591
@@
Table 4: Stability experiments. We tested the algorithm on several datasets with varied initializations and measured the standard deviation (STD) between the distinct maps. In the left table, we show a relatively small variation between solutions despite using a generative model, indicating consistent convergence. In the right table, we compared the normalized STD by dynamic range to two other XAI methods. Our approach calculates the STD over three different initializations, while in the others over three sequential layers. Notably, the layer of the classifier impacts the explanations, serving as a hyperparameter. Our method demonstrates relative stability, detailed further in the supplementary materials.
STD during Iterations
Iteration 1 10K 100K 300K
Celeb A 0.33 0.094 0.032 0.027
AFHQ 0.332 0.101 0.082 0.066
Peppers 0.312 0.099 0.035 0.052
Normalized STD by Range
Methods: Ours GC [28] I-Inf [20]
Celeb 0.024 0.098 0.072
AFHQ 0.05 0.126 0.062
Peppers 0.041 0.223 0.21
@@ 5 Discussion and Conclusion
We propose an alternative way to analyze and to visualize the reasons for classification by neural networks. It is based on decomposing the image into a part which does not contribute to the classification and to one which holds the class-related cues. This approach may not be ideal for all applications and has several limitations and drawbacks, compared to standard XAI methods: 1. The method requires training, for a specific training set and classifier (training is slow, inference is fast); 2. There is no natural ranking of the significance of pixels in the image. The amplitude of pixels in the class-distinct part can serve as a good approximation; 3. Our implementation uses GANs. The proposed concept does not rely on a GAN architecture and improvements may be achieved by diffusion-type generative models, such as [16, 36, 32, 7].
@@
Appendix
Appendix 0.A Datasets
@@
No modifications were made to the remaining datasets.
Appendix 0.B Classifiers
In our experiments three different classifiers were employed for the classification block $C$: a ResNet18-type classifier, a simple classifier described in Tab. 7, and our discriminator $D$, which possesses classification capabilities, as the output is a vector of length $c$, indicating class probabilities (details of the architecture are in Tab. 8). The test accuracy of each classifier is given in Tab. 6. We note that the discriminator exhibits the best classification performance. Thus, we selected it in the main paper to show the explanation images by the various DXAI and XAI algorithms. We conducted experiments also on ResNet18 and on the Simple classifier, showing consistently similar trends. We provide later some examples of ResNet18 results on the same datasets.
@@ Appendix 0.B Classifiers
The intentionally designed simple classifier, with lower classification ability and simple architecture, provides distinct classification results from a well-known classifier like ResNet18. This intentional dissimilarity allows for meaningful comparisons of their results and verify the connection between the classifier and its explanation, as detailed in the main paper.
Appendix 0.C Training - Additional Details
$L_{KLD}=KLD\Bigl{(}D(x^{y}),C(x^{y})\Bigl{)},$ (11)
where $KLD$ is the Kullback-Leibler distance.
$d(u,v)=\|u-v\|_{L^{1}}+\|u-v\|_{L^{2}},$
with a reconstruction loss of the form
$L_{rec}=d(x,\hat{x}^{y})+d(x,\hat{x}^{\tilde{y}}).$
The proposed loss function is
$L_{dis-rec}=d(x\odot\mathbb{I},\hat{x}^{y}\odot\mathbb{I}),$ (12)
where
and $\odot$ denotes element-wise product.
The full objective loss is:
$\begin{gathered}L_{Total}=\lambda_{adv}L_{adv}+\lambda_{KLD}L_{KLD}+\lambda_{cf}L_{class-fake}\\ +\lambda_{rec}L_{rec}+\lambda_{dr}L_{dis-rec}\end{gathered}$ (13)
In our experiments we used the following weights:
$\begin{gathered}\lambda_{adv}=2,\lambda_{KLD}=1,\lambda_{cf}=2,\\ \lambda_{rec}=4,\lambda_{dr}=4.\end{gathered}$ (14)
Appendix 0.D Inference Stage
Appendix 0.E Ablation
Figure 11: Agnostic and distinct components in two cases. Middle columns, the agnostic part is found through optimization. In this case, the distinct part resembles more noise and has little semantic features . Right columns, our method which provides much more structured and semantic components, containing features that represent the predicted class.
In the ablation study, we tried the naive method to obtain the class distinct component. As shortly mentioned in the main paper, apparently it’s possible to obtain it by optimization. The idea is to get the agnostic component and subtract it from the original image. The optimization is done as follows: Initiate $\psi_{Agnostic}=x$ when $x$ is an image from our data. Then, get the predicted probability of each class by entering the image into the classifier:
$p=C(\psi_{Agnostic}).$
Next, calculate the KL distance between the obtained distribution and the uniform distribution as follows:
$D_{KL}=\sum\limits_{i=0}^{c-1}\frac{1}{c}(\log\frac{1}{c}-\log p_{i}).$
Finally, compute iteratively or until convergence:
$\psi_{Agnostic}=\psi_{Agnostic}-dt\nabla_{\psi_{Agnostic}}D_{KL},$
and the distinct component will be $\psi_{Distinct}=x-\psi_{Agnostic}$. We received that the process does converge quickly to images for which the output is approximately a uniform distribution (With $D_{KL}<10^{-}6$). However, the results are not very informative, as can be seen in Fig. 11. Note that this is a highly non-convex problem where many local minima are possible. In the optimization case we get a result which is with very little semantic meaning. It is closer, in some sense, to adversarial attacks. This is in contrast to our method, where the results are based on the entire training set. The generative process learns how to produce features that, on average, will confuse the classifier and therefore are semantic in nature. We believe that this distinction may lead to fruitful future research, also in the case of robust network analysis and defense.
@@ Appendix 0.E Ablation
In addition, we investigated two more key factors. Firstly, we examined the impact of the number of branches. While decomposing the classified image into several images, we primarily focus on solving only two images—distinct and agnostic. One might question why not use only two branches for the solution, since two branches can be trained to achieve a similar solution.
@@
We show that when branches are used for the agnostic part results are better both in terms of reconstruction and of the generators’ ability to produce images that explain the classifier. For instance, as demonstrated in Fig. 12, PSNR decreases when using only two branches. Additionally, the loss $L_{class-fake}$, representing the generators’ ability to produce meaningful images of a specific class according to the given classifier, is higher for the entire training with two branches. In other words, the classifier interprets the images less accurately as the desired class, making the images less reliable. The additional generators provide better generation capacity and can be trained in an easier and more stable manner.
(a) AFHQ - Class-fake loss
(b) AFHQ - PSNR
(c) CelebA - Class-fake loss
(d) CelebA - PSNR
Figure 12: AFHQ & CelebA ablation - need of more than 2 branches. The run of the experiment was stopped after 150K iterations, after the tendency was understood.
@@
Appendix 0.F Stability Experiments Details
As described in the main paper, we utilize image decomposition to address various issues encountered in existing XAI methods. To ensure adaptability across different classifiers and data types, we employ generative models for generating image decompositions. However, we are aware that training using the gradient descent method does not guarantee convergence, raising concerns about the consistency of solutions across different model initializations.
Figure 13: Instability observed in Grad-CAM example with sequential layers (from Fig. 2 in the main paper).
To mitigate these concerns, we conducted experiments (Table 4) to examine whether the explanations provided by the network for classifications (specifically, ResNet18 in our experiment) become increasingly consistent throughout training, as indicated by the standard deviation of the explanations for different initializations.
Additionally, we compared this standard deviation with that of two other algorithms (Grad-CAM[28] and Internal Influence[20]), which also yield correct but potentially different solutions based on different choices of hyperparameters (in our case, the activation layer of the classifier we aim to explain). We conducted these experiments with three different initializations and three sequential layers to calculate the standard deviation.
Since the values of the pixels provided by the solutions vary between different methods, for example, in our method, the range of values is from minus 1 to 1, compared to the other two methods where the values are positive and theoretically between 0 and 1. Therefore, we normalized the standard deviation of each image within the dynamic range of the solutions, which is approximately the maximum value obtained from the three different solutions minus the minimum value. This normalization allows for a fair evaluation of quality.
Our experiments demonstrated that the standard deviation decreases with training, indicating convergence toward a consistent solution. Moreover, our standard deviation is relatively small compared to the alternatives.
Appendix 0.G More results